OpenStudy (anonymous):

how do i factor 2x^2-5x+3?

OpenStudy (anonymous):

are you familiar with ,factoring by grouping?

OpenStudy (anonymous):

yes but there are only 3 numbers.....

OpenStudy (anonymous):

well, look at the middle number, can you seperate that into 2 numbers that would allow you to factor by grouping?

OpenStudy (anonymous):

examples -5x=-2x+-3x -5x=-4x+-1x -5x=-6x+ 1x -5x= 6x + -1x

OpenStudy (anonymous):

(2x )(x ) and it has to be 3 and 1 if it's an easy factorization

OpenStudy (anonymous):

thanks also how do you factor 2x^2+7x-15

OpenStudy (anonymous):

You can use this formula: \[x _{1,2} = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] where a = 2, b = -5, c =3

OpenStudy (anonymous):

The value of a,b and c, you can get from the equation \[ax ^{2}+bx+c\] You can solve the various forms of equations with ease.

OpenStudy (anonymous):

This is just another of same (2x )(x ) and your 15 breaks out into 5 and 3

OpenStudy (anonymous):

ya but they dont add to 7 right?......

OpenStudy (anonymous):

depends on the sign and where you put each one....